the question is write and solve an equation to find the length and with of the rectangle whose width os 3 less than the length. the area is 4cm
Found 2 solutions by oberobic, solver91311:
Answer by oberobic(2304)

The area of a rectangle = length * width.
A = L * W
A = 4 (given)
W = L -3 (given)
or
L = W + 3
.
Substitute what we know now.
.
A = L*W
4 = L* (L-3)
4 = L^2 -3L
Subtract 4 from both sides
.
0 = L^2 -3L -4
L^2 -3L -4 = 0
.
Factor
(L -4)(L +1) = 0
.
L = 4 or L = -1
Well, we cannot have negative length, so L = 4.
.
Looking back, the only possible value for W = 1.
.
Done.

Answer by solver91311(24713)



Let represent the width of the rectangle (note spelling of width, not "with"). Then must represent the length. Since the area is the length times the width, represents the area. You state that the area is 4 cm but that is clearly wrong. Area is measured in square units, not linear units, hence the area must be 4 square centimeters.



And in standard form:



Just solve for , discard the negative root (since you are trying to determine a postive measure of length), and then calculate

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism